find fourier integral for the following f(x)= e-x x>0


1
Expert's answer
2022-03-09T14:54:39-0500

Solution


Given that


"f(x)=e^{-x}" here "x>0"


The Fourier expansion is defined as


"f\left( x \right) = \sum\limits_0^\infty {\left( {A\left( \alpha \right)\cos \left( {\alpha x} \right) + B\left( \alpha \right)\sin \left( {\alpha x} \right)} \right)} \"


Here,


"A\left( \alpha \right) = \frac{1}{\pi }\int\limits_{ - \infty }^\infty {f\left( u \right)\,\cos \left( {\alpha u} \right)\,du} \"

and

"B\left( \alpha \right) = \frac{1}{\pi }\int\limits_{ - \infty }^\infty {f\left( u \right)\,\sin \left( {\alpha u} \right)\,du} \"


Now, Integrating by parts, we getΒ 


"A\left( \alpha \right) = \frac{1}{{{\alpha ^2} + 1}}\" and "B\left( \alpha \right) = \frac{\alpha}{{{\alpha ^2} + 1}}\"


Hence


"\begin{array}{l}
f\left( x \right) = {e^{ - x}} = \sum\limits_0^\infty {\left( {A\left( \alpha \right)\cos \left( {\alpha x} \right) + B\left( \alpha \right)\sin \left( {\alpha x} \right)} \right)} \\
{e^{ - x}} = \sum\limits_0^\infty {\left( {\frac{1}{{{\alpha ^2} + 1}}\cos \left( {\alpha x} \right) + \frac{\alpha }{{{\alpha ^2} + 1}}\sin \left( {\alpha x} \right)} \right)} \\
{e^{ - x}} = \frac{1}{{{\alpha ^2} + 1}}\sum\limits_0^\infty {\left( {\cos \left( {\alpha x} \right) + \alpha \sin \left( {\alpha x} \right)} \right)} 
\end{array}\"



Hence, the Fourier expansion of "f(x)=e^{-x}" is

"{e^{ - x}} = \frac{1}{{{\alpha ^2} + 1}}\sum\limits_0^\infty {\left( {\cos \left( {\alpha x} \right) + \alpha \sin \left( {\alpha x} \right)} \right)} \"